1. There is a fruit seller who has 30 Kgs of apples to be sold and he wants to fix a price so that all the apples are sold. There are three customers in the market and their individual demand functions are given below:
D1=25-.05P
D2=20-.025P
D3=15-.075P
Where D is the demand and P is the price
Determine :
 Market demand equation for the fruit seller
 Price at which he can sell all the apples
 Individual demands of each of the three customers
1
Expert's answer
2015-09-24T13:08:20-0400
Market demand (D) equation for the fruit seller:
D=D1+D2+D3=25-0.05*P+20-0.025*P+15-0.075*P => D=60-0.15*P

Price at which apples will be sold:
Total quantity of apples = 30 => 30=60-0.15*P => P=200 (for price equal to 200 all the apples will be sold).

Individeal demands of each of three customers:
D1=25-0.05*200=15 => D1=15
D2=20-0.025*200=15 => D2=15
D3=15-0.075*200=0 => D3=0

So, D=D1+D2+D3=15+15+0=30 => all apples are sold